Use the following general linear supply function: Qs = 40 + 6P - 8PI + 10F where Qs is the quantity supplied of the good, P is the price of the good, PI is the price of an input, and F is the number of firms producing the good. If PI = $20, F = 60, and the demand function is Qd = 600 - 6p the equilibrium price and quantity are, respectively,
Answer #1

Qs = 40 + 6P - 8PI + 10F

PI = $20 and F = 60

So, Qs = 40 + 6P - 8(20) + 10 (60)

= 40 + 6P - 160 + 600

Qs = 480 + 6P (Supply equation)

Qd = 600- 6P (Demand equation)

Now equate Qs= Qd , we get :

480 + 6P = 600 - 6P

12P = 120

P = 120/12

P = 10

Q = 480+6(10)

= 480 + 60 = 540 units

Hence, the equilibrium price is $10 and equilibrium quantity is 540 units.